$40 MC $64 $36 Demand $56 $32 ATC $28 $48 $24 $40 Cost of Webcam $20 Price of Webcam $16 $12 $24 AVC $16 $8 $4 $8 Supply $0 $25 firms 50 firms 100 firms 80 firms
Answer #1

Answer
The market is equilibrium at D=S
where
Q=400 and P=32
for a firm, production is at P=MC
where
q=8 units
number of firms =Q/q=400/8=50
there are 50 firms in the market
Option 2